3.50 g of an unknown base is dissolved in 289 mL of water and has a concentration of 0.219 M. What is the identification of the unknown base?
1
Expert's answer
2019-09-27T04:45:47-0400

n(base) = 0.289 l * 0.219 = 0.0626 mol

M(base) = 3.50g/0.0626 mol = 55.910 ~ 56 g/mol

Then we need to know valency of metal atom in base, if metal valency is 1 then:

Molar mass of metal is 56 - 17 = 39 it’s potassium

Our base is KOH